how do you solve x-4/7=7x-5/21??
Answer by rfer(16322)

x-4/7=7x-5/21
-6x=-5/21+4/7
-6x=-5/21+12/21
-6x=7/21
x=7/21*-1/6
x=-7/126
x=-1/18